OpenStudy (mendicant_bias):

(Fourier Series) I'm having trouble understanding how to set up a given Fourier Series, example put below.

OpenStudy (jhannybean):

@ikram002p @perl

OpenStudy (mendicant_bias):

My confusion is whether L in this case (Where f(x) is a periodic function with period 2L, or -L<x<L)-whether it's pi/2 or not. half-range series tend to behave "oddly" and have a slightly different setup than normal full-range series-I understand this, but I don't totally grasp how to set them up well.

OpenStudy (mendicant_bias):

@SithsAndGiggles

OpenStudy (anonymous):

The period here is \(\pi\).

OpenStudy (mendicant_bias):

So if I were to set up a fourier cosine series where the integral's bounds run from zero to L, would the bounds be from zero to pi/2, then?

OpenStudy (anonymous):

Right, since \(2L=\pi\). Coefficients would be computed using the form \[\frac{1}{L}\int_0^{2L}\cdots\]

OpenStudy (mendicant_bias):

Oh, so it'd be from zero to pi, lol. Thank you. I've been having this issue with some other problems and want to make absolutely sure I get this, should I post them here, or do you want me to open a new question?

OpenStudy (anonymous):

Feel free. I have to go now, but I'll probably be back later.

OpenStudy (mendicant_bias):

Okay, sure thing; I'll just put it here for now. Thanks again for your help.

OpenStudy (mendicant_bias):

OS has been randomly refreshing on me, and I just lost everything I wrote... https://s3.amazonaws.com/iedu-attachments-question/875d339debeca1a951697d60d25009b2_f80f1b83e31f1e8b2386caf91f5172df.png ^For this one, I think it's pi-periodic, thus L=pi/2, and the bounds of integration for a half-series would be from 0 to 2L or 0 to pi; the coefficient out front on a_0 or a_n would be 2/L=2/(pi/2)=4/pi, is that correct? https://s3.amazonaws.com/iedu-attachments-question/46e9e4928e8049e2cbb40c6c23f1bb7e_42140d52559de40ede1cb4a1929123b8.png ^For this one, I'd think the exact same thing. And something that confuses me...is that all of my book's half-range formula-rather than putting the upper bound to 2L, it just has it to L. Granted, it multiplies the whole expression by 2.

OpenStudy (anonymous):

Yes and yes. It seems your text is assuming \(f(x)\) is symmetric across the period, which would only be okay if \(f\) is even (on the interval). For example, \(2\int_0^L\) would not work for \(f(x)=x^2\), but it would for \(f(x)=\left(x-\dfrac{\pi}{2}\right)^2\).

OpenStudy (mendicant_bias):

I get a little mixed up dealing with the notation for the coefficients because some people use a_0/2, and then use a constant *inside* that expression additionally, etc etc, if I could describe the fourier coefficients as multiplied by what constants (not writing it out, just like, "this thing multiplied by one half, regardless of formulae used"-how would I say that?

OpenStudy (anonymous):

The way I remember the formulas are by fixing the coefficient of the integrals as \(\dfrac{1}{L}\), and that the constant term of the series is \(\dfrac{a_0}{2}\).

OpenStudy (mendicant_bias):

That's a better way to remember it IMO, thanks.

OpenStudy (mendicant_bias):

I'm still a little confused about what you said here, and I have to feel it's a typo or someting: http://i.imgur.com/QcH2kH9.png If f(x)=x^2, f(x) is an even function (restating for my own sake), f(x)=f(-x), and is-at least on an interval centered on the origin, symmetric across the period, right? Wouldn't it be the exact opposite that would work, e.g. that integral would hold true for f(x)=x^2 and not f(x)=(x-pi/2)^2, with an interval centered on the origin?

OpenStudy (mendicant_bias):

My book, instead of integrating from 0 to 2L, for an even function over a symmetric interval (-L, L), multiplies the whole thing by 2, which I'm guessing is equivalent.

OpenStudy (anonymous):

What I meant was that over the interval \([0,\pi]\), the function \(f(x)=x^2\) is not symmetric about \(x=\dfrac{\pi}{2}\) (whereas \(x^2\) is symmetric across \(x=0\), but only if \(0\) is the midpoint of the interval you're considering). This means you can't use the shortcut your text mentions, that \[a_0=\frac{1}{L}\int_0^{2L}f(x)\,dx=\frac{2}{\pi}\int_0^{\pi}x^2\,dx\neq\frac{4}{\pi}\int_0^{\pi/2}x^2\,dx\] but if \(f(x)=\left(x-\dfrac{\pi}{2}\right)^2\), you can: \[a_0=\frac{2}{\pi}\int_0^{\pi}\left(x-\frac{\pi}{2}\right)^2\,dx=\frac{4}{\pi}\int_0^{\pi/2}\left(x-\frac{\pi}{2}\right)^2\,dx\]

OpenStudy (anonymous):

So while \(x^2\) IS still an even function, it's not even relative to the axis \(x=\dfrac{\pi}{2}\)

OpenStudy (anonymous):

The last thing you said is right, if \(f\) is even relative to the axis \(x=0\) (i.e. the typical symmetry seen in an even function). If \(f(x)\) is an even function defined over a symmetric interval, say \([-\pi,\pi]\), then indeed, \[a_0=\frac{1}{L}\int_{-L}^Lf(x)\,dx=\frac{2}{L}\int_0^Lf(x)\,dx\]